OpenStudy (anonymous):

sorry... i know this is a lame question but... to write an expression as a positive exponent should i just change the negative exponent to a positive one? for example.. 7^-2 would just be 7^2 right?

OpenStudy (anonymous):

You would move it to the denominator to make the exponent positive: \[a^{-n}=\frac{1}{a^n}\rightarrow 7^{-2}=\frac{1}{7^2}\]

OpenStudy (anonymous):

Oh okay so basically you just put it under 1 with the positive exponent?

OpenStudy (anonymous):

You would move it to the denominator to make the exponent positive: \[a^{-n}=\frac{1}{a^n}\rightarrow 7^{-2}=\frac{1}{7^2}\]

OpenStudy (anonymous):

yeah but sometime you end up with a reversed scenario such as: \[\frac{1}{5^{-3}}\] in this case you would move it to the numerator to make it positive \[\frac{1}{5^{-3}}\rightarrow 5^3\]

OpenStudy (anonymous):

ah cool thanks :D but i have another question XD... what if for example you have... ab^-2c or something like x^-2y^-3? i have similar problems further along but figured id ask now x]

OpenStudy (anonymous):

Is the first one \[(ab)^{-2c}, or, ab^{-2c}\]

OpenStudy (anonymous):

its \[ab ^{-2}c\] and \[x ^{-2}y ^{-3}\]sorry forgot about the equation button

OpenStudy (anonymous):

in the case of \[ab^{-2c}\] only b is raised to the negative exponent of -2c, so you would move b to the denominator\[ab^{-2c}=\frac{a}{b^{2c}}\] in the case of \[x^{-2}y^{-3}\] both x and y are raised to a negative exponent, so you would move both of them to the denominator\[x^{-2}y^{-3}=\frac{1}{x^2y^3}\]

OpenStudy (anonymous):

well the c isnt part of the b^-2 exponent its like b^-2 and c alone

OpenStudy (anonymous):

oh, sorry.... well then b will move to the denominator bc it carries a negative exponent\[ab^{-2}c=\frac{ac}{b^2}\]

OpenStudy (anonymous):

oh ok i get it now thanks for your help =]

OpenStudy (anonymous):

no problem